OpenStudy (anonymous):

PLZ HELP#VERYCONFUSED rewrite each expression in term with no power greater than 1

OpenStudy (anonymous):

\[\cos ^{3}\]

OpenStudy (anonymous):

cos^3theta

OpenStudy (anonymous):

\[\cos ^{3}\theta \]

OpenStudy (campbell_st):

it is simply \[\cos(\theta)\cos(\theta)\cos(\theta)\]

OpenStudy (anonymous):

no i have to use like identities

OpenStudy (anonymous):

ex if it is sin ^4 u do sin2)^2 and then you put the identitiy for sin^2 and then solve

OpenStudy (anonymous):

but i have no idea how to do this

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hmm I'm not familiar with any identities that decrease the power on trig functions. Are you sure it wasn't suppose to be \(cos (3\theta)\) ?

OpenStudy (anonymous):

yes i am sure but thanks

OpenStudy (campbell_st):

well start with \[\cos(\theta) (\cos^2(\theta)) = \cos(\theta)(1 - \sin^2(\theta))\]

OpenStudy (anonymous):

hmmm k

zepdrix (zepdrix):

Oh I guess we could use the Half-Angle Formulas to decrease power. I somehow forgot about those when I made my last comment :) lol

zepdrix (zepdrix):

Here's a helpful identity we can use. \[\large \color{royalblue}{\cos^2x=\frac{1+\cos2x}{2}}\]

OpenStudy (anonymous):

yeah u are right i will have to use the power reducing identities

OpenStudy (anonymous):

yeah i started by using that but i don't know what to do after that step

zepdrix (zepdrix):

\[\large \cos^3x\quad =\quad \cos x \color{royalblue}{\cos^2 x} \quad = \quad \cos x \color{royalblue}{\frac{1+\cos2x}{2}}\]

zepdrix (zepdrix):

Im not sure if that's exactly what you're looking for. But something to notice here is that we no longer even have a squared cosine, since the two cosines we're left with have different INSIDES, we can't combine them that way.

OpenStudy (anonymous):

yes i am looking for this but i think it is OK i will ask my teacher. but thanks both of u for ur help!!!